OpenStudy (anonymous):

3x+2y=-19 x-12y=19 solve using elimination

OpenStudy (whpalmer4):

the key for doing elimination is to get the equations set up so that the coefficient of one of the variables is the same magnitude but opposite in sign to that in the other equation. \[3x+2y=-19\]\[x-12y=19\] If you multiply the first equation by 6, what do you get?

OpenStudy (anonymous):

Im not shure

OpenStudy (anonymous):

Can you help me please ?

OpenStudy (whpalmer4):

Multiply each term by 6. What do you get?

OpenStudy (whpalmer4):

6*3x + 6*2y = 6*(-19)